ILPNET (Travel and subsistence claim)

The undersigned from (name in print)

(institution) attended

(name of meeting, place and date) and I am now claiming a return journey between

(my main residence)

and The cost of the return journey was

(state amount and currency)

equivalent to ecus from the budget of Additionally I am claiming days per diem at 105.40 ecus (per day).

I have read Annex II Appendix 1 of the conctract (which is reproduced below), and I certify that this claim is consistent with that agreement. Two copies of the ticket/travel agent’s certificate of costs are enclosed.

Signed Date Node manager Date

ANNEX II. Appendix 1 TRAVEL AND SUBSISTENCE

The expenses incurred in respect of the attendance of representatives of Participants at meetings convened by the Contrator in the execution of the Task shall be accepted under the contract as follows:

(a) Travel expenses Travel expenses between the usual place of work of those representatives attending the meeting and the place where the meeting is held, on production of supporting documents (certified copies of used tickets) subject to the limits set out below: If the distance by rail is greater than 400 km or the journey includes a see crossing, travel by air is acceptable. For journeys by air, the amout accepted shall not exceed the economy class air fare and the cheapest price (APEX, PEX or ”excursion”) must be sought when the difference in price to the ”Economy” class is equal to or higher than 350 ECU, or when multiple bookings are involved (excursion or ”group booking”). When a journey is undertaken by car, the amount accepted shall equal the price of a first class train ticket (including compulsory travel supplements) for the journey.

(b) Cost of shipment of equipment or unaccompaniend luggage These costs shall be accepted subject to prior approval by the Commission.

(c) Subsistence expenses The daily subsistence allowance is 105.40 ECU for each representative attending the meetings. This amount covers all expenses (hotels, meals and sundry other expenses such as taxis).